# LANCAR: Leveraging Language for Context-Aware Robot Locomotion in Unstructured Environments

### v2

@@ IV-B Evaluation Cases
In the evaluation phase, we conduct two case study experiments with increasing difficulty levels, i.e. increasing vagueness of the contextual information provided to the LLM, to examine the reasoning ability of our approach. Specifically, we evaluate the following two types of contexts (in increasing order of open-endedness):
@@ IV-B1 Low-Level Context
The contextual information provided by human observers during evaluation gives detailed qualitative descriptions of environmental properties, the same as those given in the training phase. Descriptions for all five evaluation cases we used are provided in Table II.
TABLE II: Low-Level Contextual Information for Case Study
ID Contextual Information
@@ IV-B2 High-Level Context
The contextual description provided by human observers is high-level, open-ended, vague, and descriptive of the environmental conditions, as opposed to the environmental properties. Descriptions for all five evaluation cases we used are provided in Table III.
TABLE III: High-Level Contextual Information for Case Study
ID Name Contextual Information
@@
TABLE IV: Average Episodic Reward of LANCAR and No-context RL. We perform evaluation experiments across all baselines and ablation studies over $10$ cases ($5$ low-level context cases and $5$ high-level context cases). Conclusion. ARS-based approaches achieve much higher episodic rewards than all other baselines. ARS using LANCAR embeddings for contextual information have a better performance than all other approaches in most cases.
@@ IV-D Ablation Study
We conduct a series of experiments on our approach, LANCAR, and some baseline approaches, to evaluate the effect of the usage and design of contextual information embedding strategies. We evaluate the following approaches:
@@ IV-D1 No-Context
The RL agent only uses environmental observation in their decision-making. No contextual information is used by the RL agent. The decision does not rely on the LLM output. It will be used as the baseline of the experiment.
@@ IV-D3 LANCAR (Embedding)
This is the approach we propose in this work. The LLM generates contextual embeddings by interpreting human language instruction in the way presented in Section III-C, and the RL agent incorporates contextual embeddings with environmental observation in their decision-making. The contextual embeddings are represented with a combination of multiple one-hot vectors. Each one-hot vector quantifies properties in Table I into five intervals.
@@ IV-E Results
IV-E1 Case Study: Low-Level Contextual Information
The evaluation involved testing with low-level contextual information across five different terrain cases: normal terrain (Case A), low friction (Case B), high damping (Case C), medium restitution with very high damping (Case D), and high restitution and damping with low stiffness (Case E). Generally, terrains with low friction, low stiffness, and high damping present greater challenges for RL-controlled robots, with restitution level variations increasing locomotion task uncertainty.
Table IV (Case A-E) shows the evaluation results of episodic reward over all cases using low-level contextual information. We find that LANCAR (Embedding) using an ARS backbone and Embedding method outperformed the other two approaches. LANCAR (Embedding) achieved 16.0% higher episodic rewards than LANCAR (Indexing) and 7.3% higher episodic rewards than the No-Context baseline. The LANCAR (Indexing) approach, while slightly underperforming compared to the No-Context baseline by 7.5% in episodic reward, showed variable performance across terrains, outperforming LANCAR (Embedding) in the low friction scenario (Case B) but falling behind in all other cases. This variability suggests a limited adaptation capability of the Indexing method to different terrains and contextual inputs. Notably, all methods employing an ARS backbone demonstrated better performance (approximately 17.0×17.0$\times$17.0 × higher episodic reward) than those with other backbones across all evaluation scenarios, underscoring the superior adaptability of approaches with ARS backbone in this context.
IV-E2 Case Study: High-Level Contextual Description
In our second case study, we explored high-level contextual scenarios: Moist Grassland (Case F), Snowy Mountain Road (Case G), Sunny Beach (Case H), Rainy Concrete Road (Case I), and Sunny Running Tracks (Case J). These terrains, characterized by complex combinations of surface properties like stiffness, damping, and friction, present more challenging conditions than those in the low-level context study.
Table IV (Case F-J) shows the evaluation results of episodic reward over $5$ high-level evaluation cases. Evaluation results show that LANCAR (Embedding) with an ARS backbone performed better than both LANCAR (Indexing) with 9.3% higher episodic rewards and the No-Context setups with 7.5% higher episodic rewards. LANCAR (Indexing) surpassed LANCAR (Embedding) in Case G (Snowy Mountain Road), indicating particular adaptability to low-friction conditions, but LANCAR (Embedding) excelled in all other scenarios. Approaches using the ARS backbone consistently outperformed those with different backbones across all tested terrains. Furthermore, the heightened challenge of these high-level contextual terrains accentuated the performance gap between ARS-backed methods and other approaches.
@@ IV-F Discussion
Our experiments showed that context-aware strategies, LANCAR (Indexing) and LANCAR (Embedding), consistently surpassed the baseline no-context approach in performance. LANCAR (Embedding) generally yielded superior results across diverse scenarios, showcasing its adaptability, whereas LANCAR (Indexing) excelled in specific situations. This discrepancy is attributed to the limited range of training scenarios, which may not encompass a broad spectrum of physical properties, leading to some evaluation cases falling outside the training domain. Expanding the variety of training terrains could address this issue by enhancing the model’s exposure to different environments. Notably, LANCAR (Embedding) demonstrated a more significant performance boost over no-context baselines with high-level contexts, suggesting that the LLM’s interpretation of contextual information effectively mitigates environmental ambiguity and the complexities of locomotion tasks. Additionally, methods utilizing the ARS backbone outperformed all alternatives across every scenario, affirming their superior adaptability across a wide array of evaluation conditions.
@@ V Conclusion
This paper introduces a method allowing human observers to use natural language for conveying environmental context to robots, with LLMs translating this into contextual embeddings for RL agents. These embeddings, combined with the agents’ observations, enhance navigation strategies.
Looking ahead, we aim to evolve our methodology by incorporating visual sensors and foundation models for interpreting environmental context, aiming for more accurate object captions within robot-perceived images. Another potential extension of our work is to leverage multi-modal foundation models for contextual understanding through different sensors, directly creating embeddings understandable by robots. Besides, we plan to explore mechanisms for enhancing robot adaptability across different contexts within the same episode, particularly for outdoor navigation tasks, aiming to improve the robustness and adaptiveness of real-world robot locomotion strategies.

### v3

@@ IV-B Evaluation Cases
In the evaluation phase, we conduct two case study experiments with increasing difficulty levels, i.e. increasing vagueness of the context information provided to the LLM, to examine the reasoning ability of our approach. Specifically, we evaluate the following two types of contexts (in increasing order of open-endedness):
@@ IV-B1 Low-Level Context
The context information provided by human observers during evaluation gives detailed qualitative descriptions of environmental properties, the same as those given in the training phase. Descriptions for all five evaluation cases we used are provided in Table II.
TABLE II: Low-Level Context Information for Case Study
ID Context Information
@@ IV-B2 High-Level Context
The context description provided by human observers is high-level, open-ended, vague, and descriptive of the environmental conditions, as opposed to the environmental properties. Descriptions for all five evaluation cases we used are provided in Table III.
TABLE III: High-Level Context Information for Case Study
ID Name Context Information
@@
TABLE IV: Average Episodic Reward of LANCAR and No-context RL. We perform evaluation experiments across all baselines and ablation studies over $10$ cases ($5$ low-level context cases and $5$ high-level context cases). Conclusion. ARS-based approaches achieve much higher episodic rewards than all other baselines. ARS using LANCAR embeddings for context information have a better performance than all other approaches in most cases.
@@ IV-D Ablation Study
We conduct a series of experiments on our approach, LANCAR, and some baseline approaches, to evaluate the effect of the usage and design of context information embedding strategies. We evaluate the following approaches:
@@ IV-D1 No-Context
The RL agent only uses environmental observation in their decision-making. No context information is used by the RL agent. The decision does not rely on the LLM output. It will be used as the baseline of the experiment.
@@ IV-D3 LANCAR (Embedding)
This is the approach we propose in this work. The LLM generates context embeddings by interpreting human language instruction in the way presented in Section III-C, and the RL agent incorporates context embeddings with environmental observation in their decision-making. The context embeddings are represented with a combination of multiple one-hot vectors. Each one-hot vector quantifies properties in Table I into five intervals.
@@ IV-E Results
IV-E1 Case Study: Low-Level Context Information
The evaluation involved testing with low-level context information across five different terrain cases: normal terrain (Case A), low friction (Case B), high damping (Case C), medium restitution with very high damping (Case D), and high restitution and damping with low stiffness (Case E). Generally, terrains with low friction, low stiffness, and high damping present greater challenges for RL-controlled robots, with restitution level variations increasing locomotion task uncertainty.
Table IV (Case A-E) shows the evaluation results of episodic reward over all cases using low-level context information. We find that LANCAR (Embedding) using an ARS backbone and Embedding method outperformed the other two approaches. LANCAR (Embedding) achieved 16.0% higher episodic rewards than LANCAR (Indexing) and 7.3% higher episodic rewards than the No-Context baseline. The LANCAR (Indexing) approach, while slightly underperforming compared to the No-Context baseline by 7.5% in episodic reward, showed variable performance across terrains, outperforming LANCAR (Embedding) in the low friction scenario (Case B) but falling behind in all other cases. This variability suggests a limited adaptation capability of the Indexing method to different terrains and context inputs. Notably, all methods employing an ARS backbone demonstrated better performance (approximately 17.0×17.0$\times$17.0 × higher episodic reward) than those with other backbones across all evaluation scenarios, underscoring the superior adaptability of approaches with ARS backbone in this context.
IV-E2 Case Study: High-Level Context Description
In our second case study, we explored high-level context scenarios: Moist Grassland (Case F), Snowy Mountain Road (Case G), Sunny Beach (Case H), Rainy Concrete Road (Case I), and Sunny Running Tracks (Case J). These terrains, characterized by complex combinations of surface properties like stiffness, damping, and friction, present more challenging conditions than those in the low-level context study.
Table IV (Case F-J) shows the evaluation results of episodic reward over $5$ high-level evaluation cases. Evaluation results show that LANCAR (Embedding) with an ARS backbone performed better than both LANCAR (Indexing) with 9.3% higher episodic rewards and the No-Context setups with 7.5% higher episodic rewards. LANCAR (Indexing) surpassed LANCAR (Embedding) in Case G (Snowy Mountain Road), indicating particular adaptability to low-friction conditions, but LANCAR (Embedding) excelled in all other scenarios. Approaches using the ARS backbone consistently outperformed those with different backbones across all tested terrains. Furthermore, the heightened challenge of these high-level context terrains accentuated the performance gap between ARS-backed methods and other approaches.
@@ IV-F Discussion
Our experiments showed that context-aware strategies, LANCAR (Indexing) and LANCAR (Embedding), consistently surpassed the baseline no-context approach in performance. LANCAR (Embedding) generally yielded superior results across diverse scenarios, showcasing its adaptability, whereas LANCAR (Indexing) excelled in specific situations. This discrepancy is attributed to the limited range of training scenarios, which may not encompass a broad spectrum of physical properties, leading to some evaluation cases falling outside the training domain. Expanding the variety of training terrains could address this issue by enhancing the model’s exposure to different environments. Notably, LANCAR (Embedding) demonstrated a more significant performance boost over no-context baselines with high-level contexts, suggesting that the LLM’s interpretation of context information effectively mitigates environmental ambiguity and the complexities of locomotion tasks. Additionally, methods utilizing the ARS backbone outperformed all alternatives across every scenario, affirming their superior adaptability across a wide array of evaluation conditions.
@@ V Conclusion
This paper introduces a method allowing human observers to use natural language for conveying environmental context to robots, with LLMs translating this into context embeddings for RL agents. These embeddings, combined with the agents’ observations, enhance navigation strategies.
Looking ahead, we aim to evolve our methodology by incorporating visual sensors and foundation models for interpreting environmental context, aiming for more accurate object captions within robot-perceived images. Another potential extension of our work is to leverage multi-modal foundation models for context understanding through different sensors, directly creating embeddings understandable by robots. Besides, we plan to explore mechanisms for enhancing robot adaptability across different contexts within the same episode, particularly for outdoor navigation tasks, aiming to improve the robustness and adaptiveness of real-world robot locomotion strategies.